### v1

# Sheared flow profile effects on acoustic impedance eduction in small 3D ducts

Abstract
We investigate the influence of sheared grazing flow on acoustic propagation in three-dimensional rectangular ducts. We particular focus on small ducts typical of most experimental impedance eduction facilities, for which the flow profile in the duct cross-section varies significantly. We assess the effect of simplifying this inherent two-dimensional flow profile to either a one-dimensional (2D duct) or uniform flow profile. Three flow profiles are considered, namely (i) the tensorised hyperbolic tangent, (ii) the law-of-the-wall, and (iii) a one obtained from a RANS simulation. These flow profiles are used in numerical simulations, based on the solution of the Pridmore-–Brown equation, to perform in silico impedance eduction experiments. Results show that realistic flow profiles can be approximated well in acoustics propagating waves in ducts by uniform or 2D-duct models provided the bulk Mach number is correctly accounted for, which contrasts with previous findings for more simplistic flow profiles. The key conclusion of this work is that if viscous effects are negligible and acoustic impedance is a good representation of a lined wall with grazing flow, then the simplification to a uniform flow is a reasonable approximation and traditional eduction methods are satisfactory accurate.
1 Introduction
Fan noise is a dominant source in modern turbofan aircraft, and its tonal character makes noise reduction critical for aircraft certification. Acoustic liners, typically a honeycomb core between a rigid backplate and a perforated facesheet, are the main passive treatment. Explicitly modeling the physics of a whole liner remains too expensive due to multiscale interactions — such as acoustic-induced flow and boundary layer — although small-scale high-fidelity simulations have recently become feasible [paduano_impact_2025]. Instead, liners are described by their acoustic impedance, $\tilde{Z}(\omega)=\theta+\mathrm{i}\chi$, with $\theta$ the resistance and $\chi$ the reactance. This frequency-dependent property serves as a boundary condition in acoustic models of aeroengine nacelles.
The impedance of an acoustic liner depends on liner geometry [guess1975Calculation, jones2002effects] and operating conditions such as the Sound Pressure Level (SPL) [murray2012development], grazing flow velocity amplitude [guess1975Calculation], and flow profile [kooi1981Experimental]. Today, impedance eduction is the standard experimental approach for characterizing liners with grazing flow. It infers impedance from in-duct acoustic measurements and a duct propagation model, offering simpler instrumentation, higher repeatability, and broad applicability across liner designs [watson1999validation, jing2008straightforward, elnady2009validation, ferrante2016back, bonomoComparisonSituImpedance2024].
Impedance eduction methods rely on acoustic propagation models for ducts. Common simplifications are that (i) the flow is uniform, and (ii) the Ingard–Myers Boundary Condition (IMBC) [ingard1959influence, myers1980acoustic] can be used to take into account the thin sheared boundary layer effects on acoustic propagation. These simplifications are currently under scrutiny in the scientific community. In principle, the acoustic impedance of liners should remain independent of the incident acoustic field, i.e., one can assume the locally reactive assumption. However, experiments have shown different impedance values when the wave propagation direction changes between upstream- and downstream-propagating waves, suggesting flaws in the modeling used for impedance eduction [renou2011Failure, boden2016Effect, roncen2019influence, boden2017Comparison].
Several attempts have been made to include governing equations, capturing actual liner physics, to solve for this discrepancy. For instance, Ref. [weng2018Flow] found that considering the effects of the viscosity in the acoustic propagation is unable to account for the discrepancy between upstream and downstream and collapse the educed impedances. Ref. [nark2018assessment] found that small changes to the mean flow velocity were able to reduce the gap between the upstream and downstream educed impedances, but not collapse the curves. Ref. [roncen2020WavenumberBased] showed that the difference persists even when considering a three-dimensional inviscid sheared flow profile and a no-slip condition at the lined wall, suggesting that liner physics is not fully captured. Novel boundary conditions have been introduced to better describe liner physics, but none have accurately predicted acoustic wavenumbers when the experimental configuration is changed [spillere2020Experimentally].
Recently, Ref. [bonomoUnpublished] extended the work of nayfeh1974Effect by examining how velocity profile formulations affect impedance eduction. Using the Pridmore–-Brown Equation (PBE) [pridmorebrown1958sound], wavenumbers for realistic profiles were compared against those obtained with the uniform flow assumption and the IMBC. The results showed that the IMBC [ingard1959influence, myers1980acoustic] is adequate for small ducts and low Mach numbers when using realistic profiles, but errors grow when using simplified profiles, increasing with duct width and mean flow velocity. A key limitation of Ref. [bonomoUnpublished] was the restriction to two-dimensional ducts (i.e. 1D flow profiles), leaving the role of three-dimensional effects unresolved. This raises concerns about the conclusions of Ref. [roncen2020WavenumberBased], who employed a hyperbolic tangent profile in their 3D analysis. Since the choice of velocity profile formulation strongly impacts the educed impedance, their findings may be unreliable, and a reassessment with more realistic flow representations is necessary to properly understand sheared flow effects in 3D ducts.
For clarity, we distinguish between the dimensionality of the duct, which defines the physical geometry, and that of the flow profile or solver, which determines the mathematical formulation. A three-dimensional (3D) duct refers to a rectangular geometry supporting variations in both the spanwise and wall-normal directions, for which the mean flow profile is two-dimensional. Conversely, a two-dimensional (2D) duct restricts the problem to a single wall-normal velocity profile, resulting in a one-dimensional (1D) formulation of the PBE. Throughout this paper, we refer to the 2D PBE as the partial differential form governing acoustic propagation in 3D ducts, and to the 1D Pridmore–Brown equation as the simplified ordinary differential form used in 2D ducts.
In this paper, we reassess the impact of simplifying the three-dimensional nature of acoustic wave propagation in lined rectangular ducts to either a two-dimensional domain or a uniform flow case, under conditions typical of impedance eduction facilities. Building on our recent findings on the role of axial velocity profile shape [bonomoUnpublished], we revisit earlier conclusions, including those of Ref. [roncen2020WavenumberBased], which relied on simplified flow assumptions. In this paper, we perform virtual in silico experiments using the full 2D version of the PBE as the reference model; using this, we compute the least-attenuated mode wavenumbers for different velocity profiles and then apply the straightforward impedance eduction method to quantify how uniform flow or 2D approximations may distort experimental outcomes.
This paper is organized as follows. Section 2 presents the governing equations for 3D duct acoustics with sheared grazing flow. Section 3 describes the setup for the numerical experiments, and the main results regarding the flow profile shape effects on impedance eduction in 3D ducts are presented in Section 4. Finally, the concluding remarks are presented in Section 5.
2 Governing equations
For the purpose of this study, the infinite rectangular duct depicted in Fig. 1 is considered. The duct cross-section has width $W$ and height $H$. The axial flow has flow profile $\boldsymbol{u}_{0}=U_{0}(x,y)\boldsymbol{\hat{\mathrm{e}}_{z}}$, where $\boldsymbol{\hat{\mathrm{e}}_{z}}$ is the unitary vector in the $z$ direction. Neither the flow profile nor the impedance vary in the axial direction. This implies that the flow is fully developed in the duct. The wall located at x=W/2x=-W/2 has a locally-reacting frequency-dependent impedance, $Z(\omega)$, while all the other walls are acoustically rigid.
Figure 1: Schematic duct and coordinates system adopted in this work.
For the purpose of this work, we will assume that the in-duct acoustic propagation can be described by the PBE [pridmorebrown1958sound], such that
$(\mathrm{i}\omega+\boldsymbol{u_{0}\cdot\nabla})\left(\dfrac{1}{c_{0}^{2}}(\mathrm{i}\omega+\boldsymbol{u_{0}\cdot\nabla})^{2}\tilde{p}^{\prime}-\nabla^{2}\tilde{p}^{\prime}\right)+2\dfrac{\partial}{\partial z}\big(\boldsymbol{\nabla}\tilde{p}^{\prime}\boldsymbol{\cdot\,\nabla}\!U_{0}\big)=0,$ (1)
where $\tilde{p}^{\prime}$ is the acoustic pressure with assumed monochromatic time dependence $\exp(\mathrm{i}\omega t)$, $c_{0}$ is the speed of sound, $\mathrm{i}=\sqrt{-1}$ is the complex imaginary unity and $\boldsymbol{\nabla}=(\partial/\partial x,\partial/\partial y,\partial/\partial z)$. Considering the axial invariance of the problem, we assume an axial dependence of the form $\tilde{p}^{\prime}(x,y,z)=\tilde{p}^{\prime}(x,y)\exp\{-\mathrm{i}k_{z}z\}$, where $k_{z}$ is the axial wavenumber, so Eq. (1) can be written as
$\left(\nabla_{\!\perp}^{2}+\dfrac{\omega^{2}}{c_{0}^{2}}\right)\tilde{p}^{\prime}-k_{z}\left(\dfrac{U_{0}}{\omega}\nabla_{\!\perp}^{2}-\dfrac{2}{\omega}\nabla_{\!\perp}U_{0}\cdot\nabla_{\!\perp}+\dfrac{3\omega U_{0}}{c_{0}^{2}}\right)\tilde{p}^{\prime}-k_{z}^{2}\left(1-\dfrac{3U_{0}^{2}}{c_{0}^{2}}\right)\tilde{p}^{\prime}-k_{z}^{3}\left[\dfrac{U_{0}}{\omega}\left(\dfrac{U_{0}^{2}}{c_{0}^{2}}-1\right)\right]\tilde{p}^{\prime}=0,$ (2)
where $\boldsymbol{\nabla_{\!\perp}}=(\partial/\partial x,\partial/\partial y,0)$.
The acoustic velocity $\boldsymbol{u^{\prime}}$ is then given in terms of the acoustic pressure by
$\boldsymbol{u^{\prime}}=\frac{\mathrm{i}}{\rho_{0}(\omega-U_{0}k_{z})}\boldsymbol{\nabla}\tilde{p}^{\prime}-\frac{\boldsymbol{\nabla}\tilde{p}^{\prime}\boldsymbol{\cdot\,\nabla_{\!\perp}}\!U_{0}}{\rho_{0}(\omega-U_{0}k_{z})^{2}}\boldsymbol{\hat{e}_{z}}.$ (3)
As boundary conditions, at rigid walls, the normal acoustic velocity $\boldsymbol{u}^{\prime}$ vanishes, such that
$\boldsymbol{u}^{\prime}\cdot\boldsymbol{\hat{\mathrm{n}}}=0,$ (4)
where $\boldsymbol{\hat{\mathrm{n}}}$ is a unitary normal vector pointing into the wall. Non-slip boundary condition is assumed, therefore the locally reacting impedance boundary condition can be written as
$Z=\dfrac{1}{\rho_{0}c_{0}}\dfrac{\tilde{p}^{\prime}}{\boldsymbol{u^{\prime}\cdot\hat{\mathrm{n}}}}$, (5)
where the air characteristic impedance $\rho_{0}c_{0}$ is used as a normalization factor and $\rho_{0}$ is the air density.
2.1 Uniform flow case
If a uniform flow is assumed, i.e., $U_{0}(x,y)\equiv Mc_{0}$, where $M$ is the bulk (cross-section average) Mach number, the PBE (1) reduces to the Convected Helmholtz Equation (CHE),
$\nabla^{2}\tilde{p}^{\prime}+\left(k_{0}-\mathrm{i}M\frac{\partial}{\partial z}\right)^{2}\!\!\tilde{p}^{\prime}=0.$ (6)
where $k_{0}\equiv\omega/c_{0}$ is the free-field wavenumber.
The impedance boundary condition, given by Eq. (5), is valid provided the mean flow velocity vanishes at the wall. With a uniform flow model, such as the CHE, the flow velocity adjacent to the wall is non-zero, although in reality there is a boundary layer at the wall. The refractive effects that occur within this boundary layer must be taken into account by the boundary condition. The most common approach is the IMBC [ingard1959influence, myers1980acoustic], which can be written as
$Z_{\mathrm{eff},\mathrm{IMBC}}=\frac{1}{\rho_{0}c_{0}}\dfrac{\tilde{p}^{\prime}}{\boldsymbol{\boldsymbol{u}^{\prime}}\cdot\boldsymbol{\hat{\mathrm{n}}}}=\dfrac{\omega Z}{\omega-c_{0}Mk_{z}}.$ (7)
2.2 Eigenvalue problem
We seek to describe the governing equations as a generalized eigenvalue problem. One can rewrite the PBE (Eq. (2)) as
$(\boldsymbol{\mathrm{A}}_{0}+\boldsymbol{\mathrm{A}}_{1}k_{z}+\boldsymbol{\mathrm{A}}_{2}k_{z}^{2}+\boldsymbol{\mathrm{A}}_{3}k_{z}^{3})\boldsymbol{\tilde{p}}^{\prime}=\boldsymbol{0},$ (8)
where the $\mathrm{A}_{j}$ terms involve differentiation in $x$ and $y$ and multiplication by the frequency $\omega$, by the mean flow $U_{0}$ and by its $x$- and $y$-derivatives, and $\boldsymbol{\tilde{p}}^{\prime}$ is the discretized acoustic pressure. In the present work, we follow a pseudo-spectral strategy similar to boyer2011Theoretical, where the problem is discretized by projecting it onto a Gauss–Lobatto grid using Chebyshev polynomials as basis. Finally, to solve the cubic generalized eigenvalue problem given by Eq. (8), auxiliary variables are introduced of the form $\boldsymbol{\tilde{p}}_{p}=k_{z}\boldsymbol{\tilde{p}}_{p-1}$ for $p>0$ [boyd2001Chebyshev, p. 129], and the resulting sparse linear eigenvalue problem is solved using the eigs function from Matlab.
In order to apply a lined wall boundary condition to the generalized eigenvalue problem, we rewrite Eq. (5) as
$\dfrac{\mathrm{d}\tilde{p}^{\prime}}{\mathrm{d}x}n_{x}+\dfrac{\mathrm{i}\omega}{Z}\tilde{p}^{\prime}=0,$ (9)
where $n_{x}=-1$ at x=W/2x=-W/2. For the hard wall opposite to the liner, the corresponding boundary condition is
$\dfrac{\mathrm{d}\tilde{p}^{\prime}}{\mathrm{d}x}=0.$ (10)
Similarly, for the hard walls at $y=\pm H/2$, one may write
$\dfrac{\mathrm{d}\tilde{p}^{\prime}}{\mathrm{d}y}=0.$ (11)
In this work, we employed a collocation grid with $N_{x}\times N_{y}=151\times 41$ points in the $x$ and $y$ directions, respectively. The convergence analysis was carried out independently in each direction following the same procedure as Ref. [bonomoUnpublished], yielding an estimated numerical error of $\varepsilon<$1\text{\times}{10}^{-6}$\$. It is worth noting that $N_{y}<N_{x}$ despite $H>W$, owing to the presence of only acoustically hard walls in the $y$ direction.
3 In Silico Experiments
In this section, we explore the effects of considering a 2D flow profile while solving for the acoustic propagation in a duct. We propose to revisit the work of roncen2020WavenumberBased in light of the influence of the flow profile shape formulation, by means of a numerical experiment. In this work, we consider a fully developed flow in a 3D duct.
3.1 Mean flow profile
The definition of the velocity profile for turbulent flows in a rectangular duct is a non-trivial task. The presence of secondary flows leads to the transport of flow momentum from the center of the duct to the corners, which is then pushed back to the center along the walls to preserve continuity [nakayama1983Calculation]. This creates complex patterns that make obtaining an explicit formulation for the 2D velocity profile challenging. To the best of the authors’ knowledge, no such formulation is currently available in the literature. The common simplification is to adopt a tensorisation of the 1D flow profiles in the $x$ and $y$ directions to obtain a 2D flow profile [roncen2020WavenumberBased]. In this work, we consider three flow profiles for the rectangular duct, two obtained by the traditional tensorisation procedure and one obtained with the assistance of Computational Fluid Dynamics (CFD). The three flow profile considered in this work can be see in Fig. 2.
Figure 2: Flow profiles considered in this work. (a) Tensorized law of the wall; (b) CFD evaluated, and; (c) tensorized hyperbolic tangent. Isolines represent contour axial velocity.
We use the dimensions of the Liner Impedance Test Rig from the Federal University of Santa Catarina (LITR/UFSC), which has a rectangular cross-section with a width of $W=$40\text{\,}\mathrm{mm}$\$ and height $H=$100\text{\,}\mathrm{mm}$\$, as representative of a typical liner testing facility. The frequency range is defined from 500 to $3000\,\mathrm{Hz}$, with steps of $100\,\mathrm{Hz}$. We consider a bulk Mach number of $M=0.3$.
For the tensorised flow profiles, we need formulations for 1D duct flow (channel flow). In this work, two formulations are considered. A commonly employed formulation is the hyperbolic tangent profile introduced by rienstra2008spatial, that was for instance used by roncen2020WavenumberBased to study 2D flow profile effects on impedance eduction. The hyperbolic tangent profile is given by
$M(r)=M_{c}\left[\tanh{\left(\dfrac{1-r}{\delta_{t}}\right)}+(1-\tanh{(1/\delta_{t}))\left(\dfrac{1+\tanh{(1/\delta_{t})}}{\delta_{t}}r+(1+r)\right)}(1-r)\right],$ (12)
where $M_{c}$ is the centerline Mach number, $r$ is the radial position and $\delta_{t}$ is a shape factor. In this work, we use the transformations of coordinates r(x)=|W/2x|r(x)=|W/2-x|, r(y)=|H/2y|r(y)=|H/2-y| in order to obtain the flow profiles in the $x$ and $y$ directions, respectively.
We also want to consider a more realistic representation of the boundary layer velocities profile. For this purpose, a universal law-of-the-wall is used. We consider the formulation given by van1956turbulent, such that
$U^{+}=\int_{0}^{y^{+}}\dfrac{2}{1+\sqrt{1+4\kappa^{2}{y^{+}}^{2}(1-\exp(-y^{+}/A^{+})})^{2}}\mathrm{d}y^{+}+\Pi,$ (13)
where $U^{+}\equiv U_{0}/u_{\tau}$ is the flow profile normalized by the friction velocity $u_{\tau}$, $\kappa\approx 0.42$ is the von Kármán constant, $A^{+}\approx 27$ is the Van Driest constant, and $y^{+}=\xi u_{\tau}/\nu$ is the distance from the wall, $\xi$, normalized to viscous lengths, with $\nu$ being the air kinematic viscosity. As will be discussed later, for the small ducts considered in this study, the boundary layer can extend to the entire half-duct width. To ensure that the derivative of the profile is continuous at the duct centreline, we propose adding a quadratic term to Eq. (13), denoted by $\Pi$, which is given by
$\Pi=\dfrac{2}{1+\sqrt{1+4\kappa^{2}{y^{+}_{\text{max}}}^{2}(1-\exp(-y^{+}_{\text{max}}/A^{+})})^{2}}(y^{+}_{\text{max}}-y^{+})\left(\dfrac{y^{+}}{y^{+}_{\text{max}}}\right)^{2},$ (14)
where $y^{+}_{\text{max},x}=Wu_{\tau}/\nu/2$ and $y^{+}_{\text{max},y}=Hu_{\tau}/\nu/2$ are the distances from the wall to the centreline in viscous lengths, for the $x$ and $y$ directions, respectively.
The formulations for the hyperbolic tangent and the turbulent wall law are used to generate the midspan flow profiles in the $x$ and $y$ directions ($y=0$ and $x=0$, respectively). These two 1D flow profiles are tensorised to a 2D flow profile and rescaled for the bulk Mach number by
$\dfrac{U_{0}(x,y)}{c_{0}}=M\dfrac{U_{x}(x)}{\langle U_{x}(x)\rangle}\dfrac{U_{y}(y)}{\langle U_{y}(y)\rangle},$ (15)
where the operator $\langle\cdot\rangle$ denotes averaging over the function domain. The bulk Mach number is the spatial 2D-average of the velocities profile, therefore
$M=\int_{-W/2}^{W/2}\int_{-H/2}^{H/2}\dfrac{U_{0}(x,y)}{c_{0}}\mathrm{d}y\mathrm{d}x.$ (16)
Another important quantity is the midspan average Mach number, which is often considered by research groups in in-duct propagation models [roncen2020WavenumberBased, jingInvestigationStraightforwardImpedance2015], even though it neglects the inherent two-dimensional nature of the velocities profile, and is given by
$M_{1\mathrm{D}}=\int_{-W/2}^{W/2}\dfrac{U_{0}(x,y=0)}{c_{0}}\mathrm{d}x.$ (17)
Although roncen2020WavenumberBased investigated the impact of considering either $M$ or $M_{1\mathrm{D}}$ in the wavenumbers calculations, no report is given on the impact of this choice in impedance eduction, which we will explore in this study.
The numerical simulations of the mean flow within the duct were conducted using the commercial CFD software STAR-CCM+, employing a Reynolds-Averaged Navier–Stokes (RANS) framework to resolve the time-averaged flow field. The turbulence closure was achieved using the Shear Stress Transport (SST) $k-\omega$ model developed by menter2003ten, which offers improved accuracy in capturing near-wall behavior. The simulations are carried out in a rectangular computational domain of dimensions $L_{x}\times L_{y}\times L_{z}=2.5h\times h\times 300h$, where $h=H/2=$20\text{\,}\mathrm{mm}$\$ denotes the half-height of the channel. The coordinate directions $x$, $y$, and $z$ correspond to the spanwise, wall-normal, and streamwise axes, respectively (Fig. 1). To reduce computational costs while retaining the essential flow physics, only a quarter-section of the duct is simulated. Reflection boundary conditions are imposed along the lateral and vertical walls to account for geometric symmetry and emulate the influence of the complete duct configuration.
The governing equations are discretized on a non-uniform grid consisting of $N_{x}\times N_{y}\times N_{z}=200\times 80\times 250$ cells. The streamwise discretization features anisotropic refinement, designed to capture the self-similar characteristics of the fully developed flow region. In the spanwise and wall-normal directions, grid points are clustered near the wall using a hyperbolic stretching function that ensures smooth cell-size transitions and coarsening towards the freestream. This configuration yields wall-normal spacings with $\Delta y^{+}_{\text{min}}\approx 0.58$ at the wall, and spanwise spacings between $\Delta x^{+}_{\text{min}}\approx 0.56$ and $\Delta x^{+}_{\text{max}}\approx 1.4$.
At the inflow, a prescribed velocity boundary condition is applied, while a pressure outlet is set at the downstream boundary. The freestream Mach number is $M_{\infty}=0.3$, and the friction Reynolds number is $\mathrm{Re}_{\tau}=(\delta u_{\tau})/\nu\approx 5500$.
3.2 Test Liner
[…]
Figure 3: Reference impedance considered in the numerical experiment.
4 Results and Discussion
4.1 Wavenumbers of the least-attenuated mode
First, we analyze the effects of assuming different sheared flow profiles when obtaining the wavenumbers. We compare the solutions obtained using the PBE with the tensorised hyperbolic tangent formulation (Eq. (12)), the law-of-the-wall (Eq. (13)), and CFD simulations against the solution of the convective Helmholtz equation with the IMBC, all considering the same bulk Mach number, $M$. The resulting wavenumbers are shown in Fig. 4.
Figure 4: Comparison of the wavenumbers obtained for the reference uniform impedance using different flow profile shapes. $(+)$ denotes downstream propagation, while ()(-), upstream propagation.
As previously observed for the 2D duct (with a 1D flow profile) in Ref. [bonomoUnpublished], good agreement is found between the wavenumbers obtained under the uniform flow assumption and those from the exact solution of the PBE using the law-of-the-wall profile. This suggests that, at least for plane-wavelike modes, simplifying a 3D duct to a 2D or even uniform flow representation may not lead to significant inaccuracies. This contrasts with the findings of Ref. [roncen2020WavenumberBased], who considered a less physical hyperbolic tangent flow profile. The wavenumbers obtained considering the CFD flow profile agree well with those from both the IMBC and the tensorised law-of-the-wall profile, suggesting that the latter is a good representation for the duct geometry considered. Therefore, for the sake of conciseness, we adopt the tensorised law-of-the-wall profile as a representative realistic flow profile in the remainder of this work.
4.2 On the average Mach number definition
Next, we propose an analysis similar to those of Refs. [roncen2020WavenumberBased] and [jingInvestigationStraightforwardImpedance2015] to investigate the impact of using the “correct” bulk Mach number when applying the uniform flow hypothesis or simplifying the geometry to a 2D duct. The key difference from the former lies in our inclusion of upstream-propagating waves, identified as the critical case; from the latter, our focus is on the influence of different flow profile shapes. In what follows, we compare the wavenumbers obtained in the full 3D duct with those from a 2D simplification, using two approaches: in the first (“sliced”), the midspan velocity profile is directly used, while in the second (“scaled”), this same profile is rescaled to match the bulk Mach number, $M$. Hence, the sliced profile has an average Mach number of $M_{1\mathrm{D}}$, while the scaled profile has an average Mach number of $M$ by construction. Since good agreement was previously observed between the CFD velocity profile and the law-of-the-wall representation, and because a consistent 2D reduction of the CFD profile would require a dedicated 2D duct simulation, the CFD-derived profile is omitted here for the sake of brevity. We also compare the results obtained from the IMBC when considering a uniform flow of either $M_{1\mathrm{D}}$ or $M$. The four flow profiles considered are shown in Fig. 5.
Figure 5: Comparison of the flow profiles considered on the analysis of the average flow profile effect when simplifying to a 2D duct.
Figure 6: Comparison of wavenumbers for upstream-propagating waves obtained for the reference impedance, using the fully 2D flow profile, simplified 1D profiles with and without scaling to the bulk Mach number, and the IMBC. (a,c) law-of-the-wall; (b,d) hyperbolic tangent profile.
Fig. 6 shows the wavenumbers for upstream-propagating waves and the reference impedance, considering both the tensorised law-of-the-wall and hyperbolic tangent profiles. Good agreement is observed between the full 3D solution and the 2D simplification when the 1D profile is scaled to match the bulk Mach number $M$. Moreover, the wavenumbers obtained using the IMBC with uniform flow match those from the scaled case, provided the same average Mach number is used. In contrast, the results for the hyperbolic tangent profile follow the conclusions of Ref. [roncen2020WavenumberBased]: simplifying from a 3D duct to a 2D geometry yields different wavenumbers for the same impedance, even when the flow profile is scaled. The discrepancy is even greater with the IMBC.
Figure 7: Least attenuated modes obtained for upstream propagation with the reference impedance at $3000\text{\,}\mathrm{Hz}$. (a) CHE+IMBC; (b) law-of-the-wall; (c) CFD, and; (d) hyperbolic tangent.
Fig. 7 shows the normalized acoustic pressure modal shape amplitudes obtained for the reference impedance and the different flow profiles at $3000\text{\,}\mathrm{Hz}$. One can observe that the inclusion of a non-uniform axial velocity profile introduces a vertical gradient in the modal shapes. This effect is due to the axial velocity variation across the cross-section and becomes most evident for the hyperbolic tangent profile (Fig. 7d), which exhibits a stronger velocity shear near the upper and lower walls compared to the duct centreline.
The partial conclusions of this study so far can be summarized as follows: (I) The IMBC provides a good approximation for the acoustic modes and wavenumbers in a typical impedance eduction range; and (II) the distortions caused by assuming non-realistic flow profiles may lead to distortions of the acoustic modes and errors in the wavenumbers.
4.3 Simulated impedance eduction
We now replicate another analysis from Ref. [roncen2020WavenumberBased], namely the numerical experiment involving synthesized wavenumbers, following a procedure that is more consistent with the framework developed and preliminary results found in this work. In their study, the wavenumbers obtained for the 3D duct were used in an iterative impedance eduction routine, which was then solved assuming a 2D duct geometry. However, only the sliced flow profile was considered, and no reference was made to the traditional eduction using the IMBC as an alternative, on top of considering only the simplified profile given by the hyperbolic tangent formulation.
In what follows, we apply the same procedure, but extend the analysis to include both the sliced and scaled velocity profiles for the 2D duct case, as well as the IMBC formulation using both $M$ and $M_{\mathrm{1D}}$ as input parameters. The numerical impedance eduction routine can be summarized as follows:
  1. 1.
    The PBE eigenvalue solver is used to compute the axial wavenumbers in the lined 3D duct with sheared grazing flow. A pair of axial wavenumbers is obtained, $k^{\pm}_{z,\mathrm{2D}}$, corresponding to the least attenuated modes. This step mimics a real-world impedance eduction approach where the wavenumbers are extracted from equally spaced acoustic pressure measurements using Prony-like algorithms [roncen2020WavenumberBased];
  2. 2.
    The following step depends on whether one intends to solve considering a 1D flow profile or assume a uniform flow with IMBC:
    • For the 1D case, the educed impedance is obtained by iteratively minimizing the cost function
      $\mathrm{cost~fun}=\left\|k^{\pm}_{z,\mathrm{2D}}-k^{\pm}_{z,\mathrm{1D}}\right\|,$ (18)
      which is solved independently for the upstream and downstream propagation cases;
    • For the uniform flow case with the IMBC, the classical straightforward impedance eduction routine is used, as described in Ref. [jing2008straightforward].
We consider both the tensorised law-of-the-wall and hyperbolic tangent velocity profiles for data generation in this numerical experiment. The results are shown in Fig. 8.
Figure 8: Educed impedances using the proposed numerical experiment, with wavenumbers obtained from the fully 2D (a,c) law-of-the-wall and (b,d) hyperbolic tangent flow profiles. Dashed lines correspond to impedance educed from upstream-propagating modes (downstream source), while dotted lines correspond to downstream-propagating modes (upstream source).
The analysis confirms key trends. First, consistent with Ref. [roncen2020WavenumberBased], the reduction to a 2D duct induces the classical mismatch between upstream and downstream propagation observed experimentally when the mean Mach number of the 1D profile, $M_{\mathrm{1D}}$, differs from the bulk Mach number of the 3D duct, $M$. This behavior holds for both velocity profiles considered. Second, when the 1D flow profile is scaled to match the original bulk Mach number, this discrepancy vanishes. This supports the findings of Ref. [jingInvestigationStraightforwardImpedance2015], who argued that averaging across the full cross-section yields the appropriate Mach number for eduction. Finally, impedances obtained via the traditional straightforward method, which assumes uniform flow and the Ingard–Myers boundary condition, remain a reasonable approximation, as long as the correct average Mach number is used, particularly when realistic flow profiles are employed to perform the in silico experiments.
5 Conclusions
This paper analyses the effects of two-dimensional (2D) sheared flow profiles on impedance eduction in three-dimensional (3D) ducts. The influence of adopting simplified velocity profile formulations or geometric simplifications—such as reducing the domain to 2D or assuming uniform flow with the Ingard–Myers Boundary Condition (IMBC)—was investigated. Wavenumbers obtained by solving the Pridmore–Brown equation (PBE) for different sheared flow profiles were compared to those from the Convected Helmholtz Equation (CHE) with IMBC. The results indicate that the IMBC offers a good approximation when compared to solutions based on realistic profiles, such as the universal law-of-the-wall or those derived from CFD simulations. Conversely, the widely used hyperbolic tangent profile yielded different wavenumbers. Additionally, it was shown that when reducing the domain to 2D, the average Mach number of the 1D profile must be scaled to match the bulk Mach number of the full 2D profile. A numerical experiment simulating traditional straightforward impedance eduction revealed no evidence that simplifying 3D duct acoustics to a 2D or uniform flow domain—when using realistic boundary layer representations—accounts for the directional discrepancies in educed impedance observed experimentally. This contrasts with the findings of Ref. [roncen2020WavenumberBased], who used a less realistic hyperbolic tangent profile. It is important to mention that viscous effects were neglected in this work and the investigation of the effect of viscosity in the same scenarios is suggested as future work.
Appendix A: The Goodrich (or UTAS) semi-empirical model
The Goodrich liner impedance model considered in this work is as given by yu2008validation. In this model, the liner impedance is
$Z=Z_{\text{of}}+S_{\text{r}}U_{0}+R_{\text{cm}}+\mathrm{i}\left(S_{\text{m}}U_{0}-\cot{(kh)}\right),$ (19)
where $Z_{\text{of}}$ is the perforate plate impedance, $S_{\text{r}}$ is the non-linear resistance slope, $U_{0}$ is the root-mean-squared acoustic particle velocity, which is obtained through an iterative process, where the SPL is used as an input [^6] In this work, non-linear term due to the SPL was neglected., $R_{\text{cm}}$ is the normalized grazing flow induced acoustic resistance, $S_{\text{m}}$ is the non-linear mass reactance, and $h$ is the liner cavity depth.
The impedance $Z_{\text{of}}$ is given by
$Z_{\text{of}}=\mathrm{i}\omega\dfrac{(t+\varepsilon d)}{c_{0}\sigma F\left(\frac{k_{s}d}{2}\right)},$ (20)
where $t$ is the facesheet thickness, $d$ is the perforate plate hole diameter, $\sigma$ is the percentage of open area, $F(k_{s}d/2)$ is the cross-section averaged hole velocity profile from Crandall’s solution [crandall1926theory] and
$\varepsilon d=\dfrac{d(1-0.7\sqrt{\sigma})}{1+305M^{3}}$ (21)
is the effective mass end correction [rice1971]. The cross-section-averaged hole velocity profile is defined as
$F\left(\frac{k_{s}d}{2}\right)=1-\dfrac{2J_{1}\left(\frac{k_{s}d}{2}\right)}{\frac{k_{s}d}{2}J_{0}\left(\frac{k_{s}d}{2}\right)},$ (22)
where $J_{0}$ and $J_{1}$ are zero- and first-order Bessel functions,
$k_{s}^{2}=-\mathrm{i}\dfrac{\omega\rho_{0}}{\mu}$ (23)
is the wavenumber of a viscous Stokes wave and $\mu$ is the air viscosity.
The non-linear resistance slope is given by
$S_{\text{r}}=1.336541\left(\dfrac{1-\sigma^{2}}{2c_{0}C_{d}^{2}\sigma^{2}}\right),$ (24)
where $C_{d}$ is the discharge coefficient, which for $t/d\leq 1$ is given by
$C_{d}=0.80695\sqrt{\dfrac{\sigma^{0.1}}{\exp{\left(\frac{-0.5072t}{d}\right)}}}.$ (25)
The normalized acoustic resistance under grazing flow is given by the Rice–Heidelberg derivation [heidelberg1980Experimental], such that
$R_{\text{cm}}=\dfrac{M}{\sigma\left(2+1.256\dfrac{\delta^{*}}{d}\right)},$ (26)
where $\delta^{*}$ is the flow profile boundary layer displacement thickness. Finally, the non-linear mass reactance is given by
$S_{\text{m}}=-0.0000207\dfrac{k}{\sigma^{2}}.$ (27)
Acknowledgments
On behalf of L.A. Bonomo and J.A. Cordioli, this research was partially funded by CNPq (National Council for Scientific and Technological Development). L.A. Bonomo acknowledges that this study was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior – Brasil (CAPES), Finance Code 001. E.J. Brambley gratefully acknowledges the support of the UK Engineering and Physical Sciences Research Council (EPSRC grant EP/V002929/1). The work of A. Paduano and F. Avallone is funded by the European Union (ERC, LINING, 101075903). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.